OpenStudy (anonymous):

why does (1+2t)^3dt=1/8(1+2t)^4 and not 1/4(1+2t)^4

OpenStudy (mimi_x3):

\[\int (1+2t)^3dt\]??

OpenStudy (anonymous):

Yes the answer is \[\frac{ 1 }{ 8 } (1+2t)^{4}\] but i get\[\frac{ 1 }{ 4 }(1+2t)^{4}\]

OpenStudy (mimi_x3):

Alright. Let \(u= 1+2t\)

OpenStudy (mimi_x3):

\[\int (1+2t)^{3}dt\] u = 1 + 2t => du/dt = 2 \[\int (u)^3* \frac{du}{2} => \frac{1}{2}\int(u)^3dt => \frac{1}{2}*\frac{u^4}{4} + c\]

OpenStudy (mimi_x3):

\[=> \frac{u^4}{8}+c => \frac{(1+2t)^4}{3} +c\]

OpenStudy (mimi_x3):

The last part should be: \[\frac{(1+2t)^4}{8}+c\]

OpenStudy (anonymous):

so you multiply through by the derivative of U then "undo" that outside of the integral and then solve??

OpenStudy (mimi_x3):

yee brah

OpenStudy (mimi_x3):

It's called "u-substitution"

OpenStudy (anonymous):

Thank you

OpenStudy (mimi_x3):

np brah